Grade 12th passMechanics

A pilot of mass m can bear a maximum apparent weight 7 times of mg.The aeroplane is moving in vertical circle.if the velocity of aeroplane is 210m/s while diving up from the lowest point of vertical circle, the minimum radius of vertical circle should be

Thousif MT
8 Years agoGrade 12th pass

1 Answer

Kuldip Pawar
ApprovedApproved Tutor Answer8 Years ago
m v^2/r + m g = 7 m g...(Given)v^2/r = 6 g r = (210)^2/(6×9.81) r = 749 metersSo, minimum radius of vertical circle should be 749 meters.